OpenStudy (anonymous):

determine whether the lines L1 and L2 passing through the pair of points are parallel, perpendicular, or neither. L1: (1,-3), (7,0) L2: (-7,-1), (-1,3)

OpenStudy (anonymous):

find the slope of each using \[m=\frac{y_2-y_1}{x_2-x_1}\] if you get the same number for L1 and L2 then they are parallel. if \[m_1\times m_2=-1\] i.e. if \[m_1=-\frac{1}{m_2}\] then they are perpendicular, and in neither than neither

OpenStudy (anonymous):

for instance the slope of L1 is \[m_1=\frac{0-(-3)}{-3-1}=-\frac{3}{4}\]

OpenStudy (anonymous):

other one works the same way, then compare